# BLEnD: A Benchmark for LLMs on Everyday Knowledge in Diverse Cultures and Languages

### v1

@@ 3 Construction of BLEnD
Language Coverage. We select languages with varying levels of resource availability using the metrics defined by Joshi et al. [11]. The resource availability of languages included in BLEnD is shown in Table 4 in the Appendix. Additionally, to maintain high-quality data, we ensure that at least one of our team members is a native speaker of the language and is originally from the country/region  [^2] North Korea was an exception, where we collaborated with a South Korean researcher studying North Korea. included in the dataset.
@@ 7 Limitations and Future Work
We also thank the following annotators for their invaluable help in building the dataset: Ángela Collados Ais, Kiamehr Rezaee, Nurul Ariyani, Sabrina Borrelli, Trapsilo Bumi, Helia Taheri, Chao Tan, Guanqun Cao, Dimitra Mavridou, Abderrahmane Samir Lazouni, Noufel Bouslama, Lyes Taher Khalfi, Nitumoni Neog, Bhagyashree Deka, Sikha Swarnakar, Sangeeta Neog, Nitashree Neog, Hailegnaw Tilaye, Amare Lakew, Wasihun Lakew, Yohannes Bogale, Addis Alemayehu, Yeon Su Park, Hee Su Park, Jeong Min Young, Hyewon Im, Lee Geunsoo, David Chong, Dea Adhista, Sarah Oktavianti, Muhammad Syahrul Kurniawan, Taufik Muhamad Yusup, annotators from Prolific, and all other annotators who preferred to remain unnamed.
@@ B.7 Annotation Analysis
Table 7 shows the level of agreement between the annotators, calculated by averaging the maximum votes among answers for each question in different categories across countries. Additionally, Table 8 shows the average number of answers per questions per categories across countries.
@@ C.1 Models
We use GPT-4 (gpt-4-1106-preview), GPT-3.5 (gpt-3.5-turbo-1106)  [^10] https://platform.openai.com/docs/models, Claude-3-Opus (claude-3-opus-20240229), Claude-3-Sonnet (claude-3-sonnet-20240229), Claude-3-Haiku (claude-3-haiku-20240307)  [^11] https://www.anthropic.com/api, PaLM2 (text-bison-002)  [^12] https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text, Gemini-1.0-Pro  [^13] https://ai.google.dev/gemini-api/docs/models/gemini?hl=ko, C4AI Command R+  [^14] https://docs.cohere.com/reference/about, C4AI Command R  [^15] https://huggingface.co/CohereForAI/c4ai-command-r-v01, Qwen-1.5-72B/32B/14B-Chat [4], SeaLLM-7B-v2.5 [22], Hyper CLOVA X [31], Aya-23 [3] and Aya-101 [34] for evaluation of cultural sensitivity of LLMs.

### v2

@@ 3 Construction of BLEnD
Language Coverage. We select languages with varying levels of resource availability using the metrics defined by Joshi et al. [12]. The resource availability of languages included in BLEnD is shown in Table 4 in the Appendix. Additionally, we involve at least one author who is a native speaker of the language and originally from the country/region represented in the dataset to handle the data inspection process  [^2] North Korea was an exception, where we collaborated with a South Korean researcher studying North Korean language..
@@
Table 9 in the Appendix presents the average number of ’I don’t know’ responses per question. On average, there were 1.01 out of 5 such responses per question, with a standard deviation of 0.35 (ranging from a high of 1.912 in Northern Nigeria to a low of 0.42 in South Korea). The frequency of ’I don’t know’ responses was higher in the sports and holidays/celebrations/leisure categories, likely due to questions on sports or holidays that are not widely recognized or celebrated in certain countries or regions.
@@ 7 Limitations and Future Work
We also thank the following annotators for their invaluable help in building the dataset: Ángela Collados Ais, Kiamehr Rezaee, Nurul Ariyani, Sabrina Borrelli, Trapsilo Bumi, Helia Taheri, Chao Tan, Guanqun Cao, Dimitra Mavridou, Abderrahmane Samir Lazouni, Noufel Bouslama, Lyes Taher Khalfi, Nitumoni Neog, Bhagyashree Deka, Sikha Swarnakar, Sangeeta Neog, Nitashree Neog, Hailegnaw Tilaye, Amare Lakew, Wasihun Lakew, Yohannes Bogale, Addis Alemayehu, Yeon Su Park, Hee Su Park, Jeong Min Young, Hyewon Im, Geunsoo Lee, David Chong, Dea Adhista, Sarah Oktavianti, Muhammad Syahrul Kurniawan, Taufik Muhamad Yusup, Miguel Ramirez, Thomas Welsch, annotators from Prolific, and all other annotators who preferred to remain unnamed.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Abstract and Section 1
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 7
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 7
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Section B.2
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Link provided in Abstract
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix C.1
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Figure 3(b)
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix C.1
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] URL in the Abstract.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix B.2
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix B.4, B.5, and B.6
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] See Appendix B.2
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Appendix B.2
@@ B.7 Annotation Analysis
Table 9: Average number of I don’t know for each question in different categories across countries. A value of ‘1.00’ indicates that, on average, one of the annotators failed to provide the answer to the question.
Category US GB ES MX ID CN KR DZ GR IR KP AZ JB AS NG ET
Food 0.80 0.71 0.49 0.36 0.56 0.68 0.23 1.00 0.52 0.90 1.24 0.58 1.15 0.54 1.69 0.81
Sport 1.58 1.70 1.22 1.11 1.08 0.92 0.31 1.65 1.41 1.67 1.44 1.64 1.69 1.01 2.16 1.53
Family 1.24 1.24 1.03 0.51 0.92 0.81 0.40 1.33 1.13 1.29 1.30 0.71 1.59 0.63 2.13 1.16
Education 0.92 1.02 0.83 0.39 0.48 0.37 0.24 0.82 0.58 0.52 0.51 0.37 0.69 0.25 1.42 0.61
Holidays 1.42 1.50 1.33 0.71 0.68 1.23 0.88 1.91 1.24 1.38 1.80 1.25 1.47 0.93 2.48 1.10
Work Life 0.71 1.10 0.91 0.63 0.43 0.69 0.49 0.62 1.13 1.16 1.29 0.60 1.22 0.63 1.59 0.68
Total 1.11 1.20 0.95 0.62 0.69 0.79 0.42 1.24 0.98 1.15 1.27 0.87 1.29 0.67 1.91 0.98
Table 7 shows the level of agreement between the annotators, calculated by averaging the maximum votes among answers for each question in different categories across countries. Additionally, Table 8 shows the average number of answers per questions per categories across countries. Lastly, Table 9 shows the average number of I don’t know per questions per categories across countries.
@@ C.1 Models
We use GPT-4 (gpt-4-1106-preview[24], GPT-3.5 (gpt-3.5-turbo-1106)  [^10] https://platform.openai.com/docs/models, Claude-3-Opus (claude-3-opus-20240229), Claude-3-Sonnet (claude-3-sonnet-20240229), Claude-3-Haiku (claude-3-haiku-20240307)  [^11] https://www.anthropic.com/api, Llama-3.1-70B (Llama-3.1-70B-Instruct)  [^12] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, PaLM2 (text-bison-002[2], Gemini-1.0-Pro [30], C4AI Command R+  [^13] https://docs.cohere.com/reference/about, C4AI Command R  [^14] https://huggingface.co/CohereForAI/c4ai-command-r-v01, Qwen-1.5-72B/32B/14B-Chat [5], SeaLLM-7B-v2.5 [23], HyperCLOVA X [34], Aya-23 [4] and Aya-101 [37] for evaluation of cultural sensitivity of LLMs.